# SecRepoBench: Benchmarking Code Agents for Secure Code Completion in Real-World Repositories

### v2

@@ Abstract.
This paper introduces SecRepoBench, a benchmark to evaluate code agents on secure code completion in real-world repositories. SecRepoBench has 318 code completion tasks in 27 C/C++ repositories, covering 15 CWEs. We evaluate 28 standalone LLMs and 13 code agents across 3 state-of-the-art agent frameworks using our benchmark. We find that state-of-the-art LLMs struggle with generating correct and secure code completions. However, code agents significantly outperform standalone LLMs. We show that SecRepoBench is more difficult than the prior state-of-the-art benchmark. Finally, our comprehensive analysis provides insights into potential directions for enhancing the ability of code agents to write correct and secure code in real-world repositories.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Code agents powered by Large Language Models (LLMs) have been used by tens of millions of developers (copilot-users-2025) to automate software development, boosting their productivity  (copilot-productivity; code-completion-productivity). However, they may generate code containing security vulnerabilities. Researchers have developed numerous benchmarks to evaluate the security of LLM-generated code  (pearce2022asleep; siddiq2022securityeval; hajipour2024codelmsec; wang2023enhancing; zhong2024can; fu2024constrained; peng2025cweval; yang2024seccodeplt; vero2025baxbench). While these benchmarks have provided valuable insights, they are insufficient to evaluate code agents for secure code completion tasks in real-world repositories. Existing benchmarks have the following two limitations.
Limitation: functional correctness and security tests in real-world repositories. As demonstrated by prior works (fu2024constrained; peng2025cweval), it is crucial to evaluate both the correctness and the security of LLM-generated code on the same benchmark. Without correctness evaluation, we may overestimate the security of LLM-generated code. Existing benchmarks for repository-level secure code completion tasks either have correctness tests (liang2024can), or security tests (yang2024seccodeplt), but not both on the same dataset.
Limitation: repository context. Code agents are able to navigate the entire software repository while generating code for the code completion task, e.g., by looking up function definitions, summarizing file directory structures. Existing repository-level code completion benchmarks (liang2024can; yang2024seccodeplt) provide a fixed context-retrieval method to code agents. This prevents them from evaluating various code agents that have customized context retrieval methods when exploring the software repositories.
@@
We evaluate the performance of 28 standalone LLMs as well as 13 code agents on SecRepoBench. We use 3 popular code agent frameworks (i.e., OpenHands, Aider, Claude Code) in combination with different LLMs to set up code agents. Our results show that state-of-the-art LLMs struggle to generate secure and correct code completions in the repository setting. The best LLM, GPT-5, has only 39.3% secure-pass@$1$. Compared to the previous most difficult benchmark that contains both functional correctness and security tests, BaxBench (vero2025baxbench), our benchmark SecRepoBench is even more difficult for LLMs. However, code agents significantly outperform standalone LLMs on SecRepoBench. We discover that most of the performance gains of code agents stem from generating more correct code, with relatively smaller contributions from generating more secure code. Finally, our analysis reveals that code agents still struggle with both compilation errors by hallucinating non-existent names and writing correct security checks, indicating areas for further improvement in secure and correct code completion.
@@ 2.1. Overview
SecRepoBench is a repository-level secure code completion benchmark. It contains 318 code completion tasks obtained from 27 popular GitHub C/C++ repositories covering 15 CWEs. Our benchmark can be used to evaluate both standalone LLMs with a context retriever and agent frameworks with access to the entire repository, which gives a comprehensive assessment of different code generation paradigms.
@@
Input and Output.  As shown in the left part of Figure 1, each task in SecRepoBench provides the model or code agent with two components as input: 1 A target function with an empty region to complete; 2 A repository providing context for the target function. The model or agent is then prompted to generate a code snippet that fills the empty region, which could introduce vulnerabilities if not properly implemented.
@@ 2.2. Benchmark Construction
We build SecRepoBench on top of the ARVO dataset (mei2024arvo), which contains more than 5,000 reproducible security vulnerabilities in C/C++ projects identified by Google OSS-Fuzz. We construct SecRepoBench through a systematic filtering pipeline. First, we identify security-sensitive code regions in real-world repositories. Then, we transform them into code completion tasks with corresponding developer-written correctness tests and dynamic security tests.
Secure-Sensitive Region Locating.  Following the definition of vulnerability localization (shen2021localizing), we identify the vulnerability patching region as the security-sensitive region, since the generated code in this region may reintroduce that vulnerability. We leverage real vulnerability-fixing commits to locate these security-sensitive code regions. However, as previous works (chen2023diversevul; ding2025vulnerability) have pointed out, not all changes in the fixing commits are related to the vulnerability. As a result, we adopt the labeling method from Ding et al. (ding2025vulnerability) and select tasks that have vulnerability-fixing commits modifying only one function.
To find qualified tasks, we perform multiple filtering steps on the ARVO dataset (mei2024arvo). We first deduplicate the tasks according to the fixing commit hash and filter out those that are merge commits. Next, we select tasks that patched one function in a single source code file, allowing minor changes outside the function such as comments and formatting. It should be noted that, because the security and correctness of a function depend on how that function is used by other parts of the project, SecRepoBench still tests the capability of code generation at the repository level.
@@
Correctness Test Filtering.SecRepoBench leverages developer-written unit tests from the projects to evaluate functional correctness. To guarantee the effectiveness of the unit tests, we require the test suite of each task to contain at least one passing test that is relevant to the target function given the ground truth patch. We identify those relevant tests by inserting a print statement into the target function and running the project’s test suite. The relevant unit tests are the unit tests whose output contains the print statement. We create full compilation commands for each project and run the test suite in the docker container provided by the ARVO dataset (mei2024arvo). We started our task selection from the top 40 projects by task count in the ARVO dataset. After filtering for tasks with valid unit tests, we ended up with 27 projects.
Security Test Identification.  Each task inherits a security test case (i.e., PoC) from the ARVO dataset (mei2024arvo) that causes the project to crash when the vulnerability is present. To provide a unified evaluation environment, we use the ARVO ”-fix” docker container and revert the project to the fixing commit state. We verify our evaluation setup by confirming that the project crashes with the pre-patched code and does not crash after the fixing commit when executed with the triggering PoC. This ensures that security depends solely on the code in the masked region of the fixing commit. We drop tasks that do not satisfy these criteria.
@@ 2.3. Comparison against Prior Benchmarks
We compare SecRepoBench against two state-of-the-art secure code generation benchmarks, SecCodePLT (yang2024seccodeplt) and BaxBench (vero2025baxbench). Table 1 highlights the key differences. For a fair comparison, we consider only the C/C++ secure code completion tasks of SecCodePLT.
@@ 3.1. Evaluation Setups
Metrics.  Following Chen et al. (chen2021evaluating), we use pass@$k$ to evaluate the correctness of the generated code, which measures the likelihood that at least one out of $k$ generations passes the unit tests. To evaluate both security and correctness, we adopt secure-pass@$k$ from Fu et al. (fu2024constrained), which measures the likelihood that at least one out of $k$ generations passes both unit tests and the security test. In this paper, we set $k=1$ for both metrics. In addition, to evaluate only the security of the code, we define secure percentage (secure %) as the number of tasks passing the security test over the total number of tasks, which measures the security accuracy across the benchmark.
Models and Agents.  For standalone LLM evaluation, we evaluate 28 popular large models. Based on the approach proposed by (liang2024can), we use BM25 as the context retriever, which retrieves the top 5 most relevant functions in the repository as a part of the prompt. Each function is treated as a document and tokenized using standard word-level tokenization, including punctuation. The masked target function serves as the query. For agent evaluation, we select three state-of-the-art code agent frameworks: Aider (aider2023), Openhands (openhands) and Claude Code (Anthropic2024). Code agents have access to the entire repository and can interactively explore the codebase using tools. For Aider and OpenHands, we select the top 4 performing OpenAI models and the top 2 performing Claude models based on standalone model evaluation results to serve as the underlying language models. For Claude Code, we use Claude Sonnet 4.5 Thinking as the backend model. We disable web searching tools for all agents to ensure that they rely solely on the repository context and cannot have access to external information.
Prompt.  We design a neutral prompt for all evaluations, which does not provide any security-specific instructions but simply asks the model to fill in the masked region based on the provided context. Unlike the oracle security reminder in SecCodePLT (yang2024seccodeplt) and BaxBench (vero2025baxbench) which provide CWE information to avoid task-specific vulnerabilities, our prompt reflects realistic code completion scenarios where developers do not know beforehand what vulnerabilities might be introduced. Therefore, the models must identify and prevent vulnerabilities only by understanding the task.
3.2. Results
@@
Overall Performance.  We conduct extensive evaluation on different LLMs and agent setups. Table 2 presents all results. Among standalone models, GPT-5 achieves the highest secure-pass@$1$ of 39.3% followed by OpenAI o3 at 32.4%, while the best model of the Claude series achieves 31.1%, still 8.2% behind the top OpenAI model. The lowest performing model, Llama 3.1 8B, obtained a secure-pass@$1$ of 5%, as it struggled with both correctness and security. For agents, OpenHands paired with OpenAI o3 achieves the best performance of a 53.5% secure-pass@$1$ among all setups, which shows a substantial improvement over standalone model evaluation. However, even the top setup fails to generate both secure and correct code for nearly half of the tasks, demonstrating that our repository-level benchmark is quite challenging for both standalone models and agents.
Figure 2 compares the secure-pass@$1$ of various top standalone models against BaxBench (vero2025baxbench). We observe that the performance of LLMs to generate secure and correct self-contained programs does not generalize to their performance in real-world C/C++ projects. All evaluated models show significantly lower secure-pass@$1$ on SecRepoBench compared to BaxBench. The best model GPT-5 has 53.8% secure-pass@$1$ on BaxBench, but only 39.3% on SecRepoBench. Notably, even our best agent setup (OpenHands + OpenAI o3 at 53.5%) does not surpass the standalone model performance on BaxBench (53.8%). Meanwhile, we observe that relative model ranking changes when evaluated over the two benchmarks. For example, Claude Sonnet 3.5 and OpenAI o3-mini present better results on BaxBench than GPT-4o New and DeepSeek-V3, but worse on SecRepoBench. These findings indicate that our benchmark SecRepoBench is more difficult than BaxBench.
@@
Average 12.9 17.8 12.8
Table 3. percentage improvements of different agent setups on SecRepoBench compared to standalone LLMs. On average, the increase in pass@$1$ (17.8%) is significantly higher than the increase of secure percentage (12.8%). This suggests that, compared to standalone LLMs, agents focus more on improving the functional correctness of code, but less on the underlying security issues.
@@
However, the results also reveal a key limitation of the agents, where the improvements in pass@$1$ on average are significantly higher than the improvements in secure percentage. For example, OpenHands + GPT-5 achieves a 25.5% improvement on correctness, but only has a 4.7% secure percentage increase. The agent appears to prioritize functional correctness during refinement process but lacks a reliable mechanism to identify and address security vulnerabilities it introduces during code completion, which could lead to severe security risks in real-world deployments.
In contrast, we also observe two outliers. For Aider + Claude Sonnet 4.5, it only adds 4.4% to its correctness, but improves 19.2% of the secure percentage. By manually inspecting the conversation trajectories, we found the model did check some of the vulnerabilities (e.g., initialize memory, adjust the array size, check for out-of-bound read). But it might be overly conservative and harm the functional correctness by introducing unnecessary checks or modifications, since 21% of code generated by Aider + Claude Sonnet 4.5 cannot even compile. For OpenHands + GPT-4.1, the opposite originates from GPT-4.1 struggling to effectively utilize the available tools in the OpenHands framework. Especially for the tool str_replace_editor, the model fails to use it correctly, ending up in dead loops that eventually terminate the interaction. This degrades agent performance below the baseline from some standalone models.
Reasoning vs. Non-reasoning Models.  Based on Table 2, reasoning models perform better compared to non-reasoning models in general. Among standalone LLMs, reasoning models occupy 8 out of the top 10 positions. The extended thinking process allows the models to better understand repository-level context and identify issues before generating code. But this is not always the case. For example, comparing non-reasoning Qwen3-Coder and reasoning model Qwen3 235B from the same model family, the former outperforms the latter by 6.6%. The reasoning models gpt-oss-120b and OpenAI o3-mini both provide modest results, surpassed by other more advanced models. This suggests that the quality of the reasoning mechanism and the degree to which the model is calibrated for secure coding tasks also matter.

### v3

@@ Abstract.
This paper introduces SecRepoBench, a benchmark to evaluate code agents on secure code completion in real-world repositories. SecRepoBench has 318 code completion tasks in 27 C/C++ repositories, covering 15 CWEs. We evaluate 29 standalone LLMs and 15 code agents across 3 state-of-the-art agent frameworks using our benchmark. We find that state-of-the-art LLMs struggle with generating correct and secure code completions. However, code agents significantly outperform standalone LLMs. We show that SecRepoBench is more difficult than the prior state-of-the-art benchmark. Finally, our comprehensive analysis provides insights into potential directions for enhancing the ability of code agents to write correct and secure code in real-world repositories.
Secure Code Completion, LLM Agent, Benchmarks
journalyear: 2026copyright: ccconference: 3rd International Workshop on Large Language Models For Code ; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 3rd International Workshop on Large Language Models For Code (LLM4Code ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3786181.3788703isbn: 979-8-4007-2412-1/2026/04ccs: Security and privacy Software and application security
@@ 1. Introduction
Code agents powered by Large Language Models (LLMs) have been used by tens of millions of developers (Business of Apps, 2025) to automate software development, boosting their productivity  (Eirini Kalliamvakou, GitHub Blog, 2022; Maxim Tabachnyk and Stoyan Nikolov, Google Research, 2022). However, they may generate code containing security vulnerabilities. Researchers have developed numerous benchmarks to evaluate the security of LLM-generated code  (Pearce et al., 2022; Siddiq and Santos, 2022; Hajipour et al., 2024; Wang et al., 2023; Zhong and Wang, 2024; Fu et al., 2024; Peng et al., 2025; Nie et al., 2025; Vero et al., 2025). While these benchmarks have provided valuable insights, they are insufficient to evaluate code agents for secure code completion tasks in real-world repositories. Existing benchmarks have the following two limitations.
Limitation: functional correctness and security tests in real-world repositories. As demonstrated by prior works (Fu et al., 2024; Peng et al., 2025), it is crucial to evaluate both the correctness and the security of LLM-generated code on the same benchmark. Without correctness evaluation, we may overestimate the security of LLM-generated code. Existing benchmarks for repository-level secure code completion tasks either have correctness tests (Liang et al., 2024), or security tests (Nie et al., 2025), but not both on the same dataset.
Limitation: repository context. Code agents are able to navigate the entire software repository while generating code for the code completion task, e.g., by looking up function definitions, summarizing file directory structures. Existing repository-level code completion benchmarks (Liang et al., 2024; Nie et al., 2025) provide a fixed context-retrieval method to code agents. This prevents them from evaluating various code agents that have customized context retrieval methods when exploring the software repositories.
@@
We evaluate the performance of 29 standalone LLMs as well as 15 code agents on SecRepoBench. We use 3 popular code agent frameworks (i.e., OpenHands, Aider, Claude Code) in combination with different LLMs to set up code agents. Our results show that state-of-the-art LLMs struggle to generate secure and correct code completions in the repository setting. The best LLM, GPT-5, has only 39.3% secure-pass@$1$. Compared to the previous most difficult benchmark that contains both functional correctness and security tests, BaxBench (Vero et al., 2025), our benchmark SecRepoBench is even more difficult for LLMs. However, code agents significantly outperform standalone LLMs on SecRepoBench. We discover that most of the performance gains of code agents stem from generating more correct code, with relatively smaller contributions from generating more secure code. Finally, our analysis reveals that code agents still struggle with both compilation errors by hallucinating non-existent names and writing correct security checks, indicating areas for further improvement in secure and correct code completion.
@@ 2.1. Overview
SecRepoBench [^1] Code and dataset can be found at: https://secrepobench.github.io/ is a repository-level secure code completion benchmark. It contains 318 code completion tasks obtained from 27 popular GitHub C/C++ repositories covering 15 CWEs. Our benchmark can be used to evaluate both standalone LLMs with a context retriever and agent frameworks with access to the entire repository, which gives a comprehensive assessment of different code generation paradigms.
@@
Input and Output.  As shown in the left part of Figure 1, each task in SecRepoBench provides the model or code agent with two components as input: [Uncaptioned image] A target function with an empty region to complete; [Uncaptioned image] A repository providing context for the target function. The model or agent is then prompted to generate a code snippet that fills the empty region, which could introduce vulnerabilities if not properly implemented.
@@ 2.2. Benchmark Construction
We build SecRepoBench on top of the ARVO dataset (Mei et al., 2024), which contains more than 5,000 reproducible security vulnerabilities in C/C++ projects identified by Google OSS-Fuzz. We construct SecRepoBench through a systematic filtering pipeline. First, we identify security-sensitive code regions in real-world repositories. Then, we transform them into code completion tasks with corresponding developer-written correctness tests and dynamic security tests.
Secure-Sensitive Region Locating.  Following the definition of vulnerability localization (Shen et al., 2021), we identify the vulnerability patching region as the security-sensitive region, since the generated code in this region may reintroduce that vulnerability. We leverage real vulnerability-fixing commits to locate these security-sensitive code regions. However, as previous works (Chen et al., 2023; Ding et al., 2025) have pointed out, not all changes in the fixing commits are related to the vulnerability. As a result, we adopt the labeling method from Ding et al. (Ding et al., 2025) and select tasks that have vulnerability-fixing commits modifying only one function.
To find qualified tasks, we perform multiple filtering steps on the ARVO dataset (Mei et al., 2024). We first deduplicate the tasks according to the fixing commit hash and filter out those that are merge commits. Next, we select tasks that patched one function in a single source code file, allowing minor changes outside the function such as comments and formatting. It should be noted that, because the security and correctness of a function depend on how that function is used by other parts of the project, SecRepoBench still tests the capability of code generation at the repository level.
@@
Correctness Test Filtering.SecRepoBench leverages developer-written unit tests from the projects to evaluate functional correctness. To guarantee the effectiveness of the unit tests, we require the test suite of each task to contain at least one passing test that is relevant to the target function given the ground truth patch. We identify those relevant tests by inserting a print statement into the target function and running the project’s test suite. The relevant unit tests are the unit tests whose output contains the print statement. We create full compilation commands for each project and run the test suite in the docker container provided by the ARVO dataset (Mei et al., 2024). We started our task selection from the top 40 projects by task count in the ARVO dataset. After filtering for tasks with valid unit tests, we ended up with 27 projects.
Security Test Identification.  Each task inherits a security test case (i.e., PoC) from the ARVO dataset (Mei et al., 2024) that causes the project to crash when the vulnerability is present. To provide a unified evaluation environment, we use the ARVO ”-fix” docker container and revert the project to the fixing commit state. We verify our evaluation setup by confirming that the project crashes with the pre-patched code and does not crash after the fixing commit when executed with the triggering PoC. This ensures that security depends solely on the code in the masked region of the fixing commit. We drop tasks that do not satisfy these criteria.
@@ 2.3. Comparison against Prior Benchmarks
We compare SecRepoBench against two state-of-the-art secure code generation benchmarks, SecCodePLT (Nie et al., 2025) and BaxBench (Vero et al., 2025). Table 1 highlights the key differences. For a fair comparison, we consider only the C/C++ secure code completion tasks of SecCodePLT.
@@ 3.1. Evaluation Setups
Metrics.  Following Chen et al. (Chen et al., 2021), we use pass@$k$ to evaluate the correctness of the generated code, which measures the likelihood that at least one out of $k$ generations passes the unit tests. To evaluate both security and correctness, we adopt secure-pass@$k$ from Fu et al. (Fu et al., 2024), which measures the likelihood that at least one out of $k$ generations passes both unit tests and the security test. In this paper, we set $k=1$ for both metrics. In addition, to evaluate only the security of the code, we define secure percentage (secure %) as the number of tasks passing the security test over the total number of tasks, which measures the security accuracy across the benchmark.
Models and Agents.  For standalone LLM evaluation, we evaluate 29 popular large models. Based on the approach proposed by (Liang et al., 2024), we use BM25 as the context retriever, which retrieves the top 5 most relevant functions in the repository as a part of the prompt. Each function is treated as a document and tokenized using standard word-level tokenization, including punctuation. The masked target function serves as the query. For agent evaluation, we select three state-of-the-art code agent frameworks: Aider (Gauthier, 2023), Openhands (Wang et al., 2024) and Claude Code (Anthropic, 2025). Code agents have access to the entire repository and can interactively explore the codebase using tools. For Aider and OpenHands, we select the top 4 performing OpenAI models, the top 2 performing Claude models, and the top 1 performing Gemini model based on standalone model evaluation results to serve as the underlying language models. For Claude Code, we use Claude Sonnet 4.5 Thinking as the backend model. We disable web searching tools for all agents to ensure that they rely solely on the repository context and cannot have access to external information.
@@
Prompt.  We design a neutral prompt for all evaluations, which does not provide any security-specific instructions but simply asks the model to fill in the masked region based on the provided context. Unlike the oracle security reminder in SecCodePLT (Nie et al., 2025) and BaxBench (Vero et al., 2025) which provide CWE information to avoid task-specific vulnerabilities, our prompt reflects realistic code completion scenarios where developers do not know beforehand what vulnerabilities might be introduced. Therefore, the models must identify and prevent vulnerabilities only by understanding the task.
3.2. Results
Overall Performance.  We conduct extensive evaluation on different LLMs and agent setups. Table 2 presents all results. Among standalone models, GPT-5 achieves the highest secure-pass@$1$ of 39.3% followed by OpenAI o3 at 32.4%, while the best model of the Claude series achieves 31.1%, still 8.2% behind the top OpenAI model. The best Gemini model achieves an even lower score of 27.7%. The lowest performing model, Llama 3.1 8B, obtained a secure-pass@$1$ of 5%, as it struggled with both correctness and security. For agents, OpenHands paired with OpenAI o3 achieves the best performance of a 53.5% secure-pass@$1$ among all setups, which shows a substantial improvement over standalone model evaluation. However, even the top setup fails to generate both secure and correct code for nearly half of the tasks, demonstrating that our repository-level benchmark is quite challenging for both standalone models and agents.
Figure 2 compares the secure-pass@$1$ of various top standalone models against BaxBench (Vero et al., 2025). We observe that the performance of LLMs to generate secure and correct self-contained programs does not generalize to their performance in real-world C/C++ projects. All evaluated models show significantly lower secure-pass@$1$ on SecRepoBench compared to BaxBench. The best model GPT-5 has 53.8% secure-pass@$1$ on BaxBench, but only 39.3% on SecRepoBench. Notably, even our best agent setup (OpenHands + OpenAI o3 at 53.5%) does not surpass the standalone model performance on BaxBench (53.8%). Meanwhile, we observe that relative model ranking changes when evaluated over the two benchmarks. For example, Claude Sonnet 3.5 and OpenAI o3-mini present better results on BaxBench than GPT-4o New and DeepSeek-V3, but worse on SecRepoBench. These findings indicate that our benchmark SecRepoBench is more difficult than BaxBench.
@@ 3.2. Results
Average 13.7 18.3 14.5
Table 3. Percentage improvements of different agent setups on SecRepoBench compared to standalone LLMs. On average, the increase in pass@$1$ (18.3%) is significantly higher than the increase of secure percentage (14.5%). This suggests that, compared to standalone LLMs, agents focus more on improving the functional correctness of code, but less on the underlying security issues.
@@
However, the results also reveal a key limitation of the agents, where the improvements in pass@$1$ on average are significantly higher than the improvements in secure percentage. For example, OpenHands + GPT-5 achieves a 25.5% improvement in correctness, but only has a 4.7% secure percentage increase. The agent appears to prioritize functional correctness during refinement process but lacks a reliable mechanism to identify and address security vulnerabilities it introduces during code completion, which could lead to severe security risks in real-world deployments.
In contrast, we also observe four outliers. For Aider + Claude Sonnet 4.5, it only adds 4.4% to its correctness, but improves 19.2% of the secure percentage. By manually inspecting the conversation trajectories, we found the model did check some of the vulnerabilities (e.g., initialize memory, adjust the array size, check for out-of-bound read). But it might be overly conservative and harm the functional correctness by introducing unnecessary checks or modifications, since 21% of code generated by Aider + Claude Sonnet 4.5 cannot even compile. For OpenHands + GPT-4.1, the opposite originates from GPT-4.1 struggling to effectively utilize the available tools in the OpenHands framework. Especially for the tool str_replace_editor, the model fails to use it correctly, ending up in dead loops that eventually terminate the interaction. This degrades agent performance below the baseline from some standalone models. For Gemini 3 Pro, both agent combinations have a smaller improvement in correctness. This is mainly because the standalone Gemini 3 Pro baseline has a much lower secure percentage comparing to other models in the same tier.
Reasoning vs. Non-reasoning Models.  Based on Table 2, reasoning models perform better compared to non-reasoning models in general. Among standalone LLMs, reasoning models occupy 9 out of the top 10 positions. The extended thinking process allows the models to better understand repository-level context and identify issues before generating code. But this is not always the case. For example, comparing non-reasoning Qwen3-Coder and reasoning model Qwen3 235B from the same model family, the former outperforms the latter by 6.6%. The reasoning models gpt-oss-120b and OpenAI o3-mini both provide modest results, surpassed by other more advanced models. This suggests that the quality of the reasoning mechanism and the degree to which the model is calibrated for secure coding tasks also matter.